OpenStudy (anonymous):

What are the possible positive, negative, or complex possibilities of 4x^3-9x^2-7x-6

OpenStudy (anonymous):

Do you know the rational root theorem?

OpenStudy (anonymous):

No I'm new to this. I just want the answer so I can do the others

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

then you can solve that and find complex roots.

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

\( f(3)=0 \)

OpenStudy (anonymous):

So 0 for positive?

OpenStudy (anonymous):

nope, what's between in the brackets of f(x) is what you insert into the function, therefore 3 is a positive, real root.

OpenStudy (anonymous):

Yes but I need the positive, negative, and compleex possible roots

OpenStudy (anonymous):

then you will get as a quadratic polynomial \( f(x) = 4x^2+3x+2\) there you only have complex roots.

OpenStudy (anonymous):

No negatives?

OpenStudy (anonymous):

hmm yes negative complex roots, but no real negative roots

OpenStudy (anonymous):

for the complex roots you will get: \( \\ \large x_1 = \frac{1}{8}(-3-i\sqrt{23})\)

OpenStudy (anonymous):

and therefore also it's complex conjugate.

OpenStudy (anonymous):

compare the given equation with \[a {x ^{2}} + b y^{2}+c\]

OpenStudy (anonymous):

How many total zeros

OpenStudy (anonymous):

three, it is a polynomial of third degree.

OpenStudy (anonymous):

therefore three roots in \(\mathbb{C}\)

OpenStudy (anonymous):

So there's 3 complex

OpenStudy (anonymous):

nope, the complex numbers include the reals, it's a higher set of numbers. One root \( \in \mathbb{R}\) and two roots \( \in \mathbb{C}\)

OpenStudy (anonymous):

Nevermind -____-

OpenStudy (anonymous):

where's the problem?

OpenStudy (anonymous):

now substitute the values of a,b,c in th equation below\[[-b+\sqrt{(b ^{2}-4a.c)} ]\div2.a\]

OpenStudy (anonymous):

well done @mbibin91, by the way, if you don't already know, if you want a nice clean multiplication dot, use \cdot in LaTeX

OpenStudy (anonymous):

There's 1 positive 2 or 0 negative and 2 or 0 complex

OpenStudy (anonymous):

@Spacelimbus k :)

OpenStudy (anonymous):

1 positive, 0 negative (in reals) and 2 complex